A chemist mixes an 11% acetic solution with a 6% acetic acid solution. How many mililiters of each should the chemist use to make 600 ml of a solution that is 8% acetic acid

Found 2 solutions by josgarithmetic, richwmiller:
Answer by josgarithmetic(39838)
Answer by richwmiller(17219)

a+b=600,
0.11*a+0.06*b=0.08*600
a=600-b
0.11*(600-b)+0.06*b=48
66-0.11b+0.06*b=48
-0.05*b=-18
b=360
a=600-b
a=240 mL at 11% b=360 mL at 6%
check
0.11*240+0.06*360=0.08*600
26.4+21.6=48
48=48
ok